OpenStudy (anonymous):

Solve the proportion using number scene 2/8 = x/48 MEDAL A. 6 B. 12 C. 16 D. 8

OpenStudy (anonymous):

48 Anna5656665 is that your reall picture

OpenStudy (anonymous):

My profile pic yeah. So my answer is 48???

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what provi it humphriescody

OpenStudy (anonymous):

I put up choices

OpenStudy (texaschic101):

I can solve the proportion, I just do not know what number scene is...lol 2/8 = x/48 -- cross multiply 8x = (2)(48) 8x = 96 x = 96/8 x = 12

OpenStudy (anonymous):

Thx! <3

OpenStudy (anonymous):

12

OpenStudy (anonymous):

oh my bad i did not meen to put that one in this windo